A point charge is moving with speed 2 × 107 m/s parallel to the x axis along strait line at y=2 m. At t = 0, the charge is at x = 0 m. The magnitude of the magnetic field at x = 4 m is B0 at the origin. The magnitude of the magnetic field at at origin when t = 0.15 μs is
Answer #1

given that

r1 = 4 m

r2 = 2m

now as we know field due to moving charge

B = ( uo / 4 pi) 4 ( v × r) / r^3

so B is inversely proportional to square of the distance

so

B2 /B1 = (d1 / d2) ^2

B2 = Bo ( 4 / 2)^2

B2 = 4 Bo

======

comment in case any doubt, will reply for sure.. goodluck